If you wish to estimate a population mean with a sampling distribution error SE= 0.33 using a 95% confidence interval and you know from prior sampling that is approximately equal to 7.6, how many observations would have to be included in your sample? The number of observations that would have to be included in your sample is (Round up to the nearest observation.)
Answer #1

SE = 0.33

Varience = 7.6 , standard deviation = sigma = sqrt(7.6) = 2.76

Standard error = SE = sigma / sqrt ( n )

0.33 = 2.76 / sqrt (n)

Sqrt ( n ) = 2.76 / 0.33

n = ( 2.76 / 0.33 ) 2 = 69.79 = 70